### v1

# ATHAR: A High-Quality and Diverse Dataset for Classical Arabic to English Translation

Abstract
Classical Arabic represents a significant era, encompassing the golden age of Arab culture, philosophy, and scientific literature. With a broad consensus on the importance of translating these literatures to enrich knowledge dissemination across communities, the advent of large language models (LLMs) and translation systems offers promising tools to facilitate this goal. However, we have identified a scarcity of translation datasets in Classical Arabic, which are often limited in scope and topics, hindering the development of high-quality translation systems. In response, we present the ATHAR dataset, comprising 66,000 high-quality Classical Arabic to English translation samples that cover a wide array of subjects including science, culture, and philosophy. Furthermore, we assess the performance of current state-of-the-art LLMs under various settings, concluding that there is a need for such datasets in current systems. Our findings highlight how models can benefit from fine-tuning or incorporating this dataset into their pretraining pipelines. The dataset is publicly available on the HuggingFace Data Hub at https://huggingface.co/datasets/mohamed-khalil/ATHAR.
1 Introduction
Classical Arabic is the foundation of Arabic linguistic theory and is well comprehended by educated Arabic readers. It significantly differs from Modern Standard Arabic (MSA) it is also called (Arangiyya), which is more simplified in terms of its vocabulary, syntax, morphology, phraseology, and semantics.
Classical Arabic poses unique challenges for accurate translation into English. Unlike MSA, which dominates formal speeches, news channels, and modern literary works, and urban dialects prevalent on social media platforms, Classical Arabic is less commonly used today. Yet, it remains vital, present in many historical documents, books, and literary texts rich with knowledge from the Arab and Muslim golden ages, all awaiting translation and broader exposure.
Current translation systems, including Google Translate and large language models like ChatGPT and Llama, struggle with Classical Arabic, often neglecting it in favour of MSA and urban dialects during dataset creation for machine translation.
This work introduces the ATHAR dataset, a translation resource from Classical Arabic to English. “ATHAR” “\<أثر>” means "legacy" or "ancient work." It represents the literary and cultural heritage and underscores the dataset’s role in illuminating classical Arabic texts, emphasizing their importance in preserving and conveying this heritage. The ATHAR dataset aims to address the representativeness and quality limitations of previous datasets.
This work is organised as follows: Section 2 explores the challenges faced by previous researchers in translating Classical Arabic and details how the ATHAR dataset addresses these challenges. Section 3 elaborates on the methodologies used to create the ATHAR dataset, including steps for data collection, cleaning, and preprocessing to ensure the quality and reliability of the data. In Section 4 we conduct experiments to assess the performance of state-of-the-art LLMs on the ATHAR dataset across various settings such as zero-shot, few-shot, and fine-tuning scenarios. The paper concludes with Section 5, highlighting the importance of the ATHAR dataset in developing culturally and linguistically authentic Arabic language models and advancing Arabic natural language processing.
2 Related Work
The notable gap in datasets for Classical Arabic has led to several efforts to gather more resources for Arabic Natural Language Processing (NLP). Prominent among these are the Tanzil and Authentic Hadith datasets, which draw from religious texts. The Tanzil dataset offers translations of the Quran in over 40 languages, including Arabic to English, and is hosted on Tanzil.net and the OPUS database Tiedemann (2012). The Authentic Hadith dataset provides translations of the sayings and practices of the Prophet Muhammad, known for its authenticity and rigorous translation process Altammami et al. (2020). While these datasets are rich, they mainly focus on religious content and don’t fully represent the diverse genres of classical Arabic literature. Additionally, the Poem Comprehensive Dataset (PCD) Yousef et al. (2019) provides a dataset focused on Classical Arabic poetry. While this dataset is a valuable resource, it encompasses a limited range of thematic areas.
In contrast, there are numerous datasets for Modern Arabic that include a rich and diverse context, such as the OPUS-100 dataset Zhang et al. (2020), the MultiUN dataset Eisele and Chen (2010), and the IWSLT2017 dataset Cettolo et al. (2017). However, Modern Arabic differs significantly from Classical Arabic in its vocabulary, syntax, and stylistic features, which are not well-represented in these contemporary datasets.
Additionally, significant efforts like those by Alrabiah et al. (2014) have focused on Arabic historical linguistics, producing datasets that explore the evolution and contexts of the Arabic language. Although these datasets are not directly applicable in practical translation tasks due to their lack of translations into other languages, they offer invaluable resources for pretraining LLMs with the knowledge necessary to distinguish between Classical and Modern Arabic. Moreover, the initiative by Aloui et al. (2024) introduced a corpus of 101 billion Arabic words, crucial for developing LLMs targeted at the Semitic Arabic language. This extensive corpus, predominantly in Modern Arabic with some Classical content, could help LLMs understand Classical Arabic, particularly when combined with smaller, specialized downstream translation datasets.
ATHAR dataset aims to address the representativeness issues in previous classical Arabic datasets by compiling sentences from various contexts and historical periods on topics like science, medicine, philosophy, and culture. This dataset will help fill the gaps in classical Arabic resources and provide a more comprehensive foundation for developing effective translation models.
3 ATHAR Dataset
This section outlines the development of the ATHAR dataset. We start by identifying the sources from which the data was collected. Subsequently, we detail the processing steps implemented to ensure the dataset’s high quality. Additionally, we compare ATHAR to previous classical Arabic datasets and well-known modern Arabic datasets.
3.1 Data Collection
The ATHAR dataset comprises 66K sentences extracted from seminal Classical Arabic texts, expertly translated into English. These texts are cornerstone works spanning diverse genres and historical periods, providing profound insights into Islamic and world history, philosophy, science, medicine, and culture. This rich literary heritage is showcased in our dataset, sourced from the Rasaif website at https://rasaif.com/ [^1] At the time of data collection and publication of this work, there were no restrictions on scraping and using resources from this website.. The specific books included in our dataset are detailed below:
    تاريخ الطبري> (History of al-Tabari): Authored by the Persian historian Muhammad ibn Jarir al-Tabari, this comprehensive history covers the world from creation to 915 AD and is a pivotal source for early Islamic history.
    تُحفة النُّظار - الرحلة> (The Travels of Ibn Battuta): This travelogue by Ibn Battuta documents his 14th-century travels across the Islamic world and beyond, offering insights into the era’s geography, cultures, and politics.
    مقدمة ابن خلدون> (The Muqaddimah of Ibn Khaldun): A pioneering work in historiography and sociology by Ibn Khaldun that analyzes the rise and fall of empires and the dynamics of society, completed in 1377 AD.
    الاموال> (The Book of Revenue): Abu Ubayd al-Qasim ibn Sallam’s exploration of Islamic economic principles, taxation, and public finance, written around 837 AD.
    العقد الفريد> (The Unique Necklace): A collection by Ibn Abd Rabbih encompassing Arabic prose and poetry on literature, history, and moral teachings, finalized in the early 10th century.
    المناظر> (The Optics): Ibn al-Haytham’s foundational text in optics that discusses the nature of light and vision, pivotal to the development of the scientific method, completed around 1021 AD.
    النوادر السلطانية و المحاسن اليوسفية> (The Sultan’s Anecdotes and Yusuf’s Merits): A biographical account of Sultan Salah al-Din by Ibn Shaddad, emphasizing his military and administrative achievements, written in the late 12th century.
    التصريف لمن عجز عن التأليف> (The Method of Healing): Al-Zahrawi’s extensive medical encyclopedia that details surgical procedures and treatments, influential in medieval European medicine, completed around 1000 AD.
    نشوار المحاضرة و أخبار المذاكرة> (Anecdotes of the Session and Stories of Recollection): Al-Tanukhi’s compilation of stories and historical accounts that offer a window into the social and cultural life of the Islamic world, from the 10th century.
    القانون في الطب> (The Canon of Medicine): Ibn Sina’s comprehensive medical guide that became a foundational text in Islamic and European medicine, covering various health aspects, finalized around 1025 AD.
    الاعتبار> (The Book of Reflection): Usama ibn Munqidh’s autobiographical reflections on the Crusades and life in the medieval Islamic world, completed in the mid-12th century.
    الرسالة> (The Epistle): Imam al-Shafi’i’s seminal work on Islamic jurisprudence, outlining legal theory and methodology, completed around 820 AD.
    البخلاء> (The Book of Misers): Al-Jahiz’s satirical critique of miserliness, offering insights into the social and economic conditions of his time, completed in the 9th century.
    نَهْجُ البَلاغَةِ> (The Path of Eloquence): A compilation of sermons, letters, and sayings attributed to Ali ibn Abi Talib, noted for its literary quality and philosophical depth, assembled in the 10th century.
    فتوح الشام> (Fattouh Al-Sham): Al-Waqidi’s detailed historical account of the Islamic conquests in Syria, documenting battles and strategies of the era, completed in the 9th century.
    الأخلاق والسير> (Ethics and Conduct): Ibn Miskawayh’s treatise on ethics and morality from an Islamic philosophical perspective, penned in the 11th century.
    حي بن يقظان> (Hayy ibn Yaqdhan): Ibn Tufail’s philosophical novel about a boy’s journey to intellectual and spiritual enlightenment on a deserted island, completed in the 12th century.
    الاصنام> (The Book of Idols): Hisham ibn al-Kalbi’s detailed account of pre-Islamic Arabian deities and their worship rituals, written in the 9th century.
3.2 Preprocessing
To prepare the dataset for use in machine translation models, several preprocessing steps were undertaken:
Cleaning the Data:
During the initial stages of the ATHAR dataset collection process, the primary challenge we encountered involved entries where Arabic and English texts were flipped within HTML class labels. For further details on this issue, see Appendix A. To address this, we implemented a simple rule-based technique that identifies the language of the text based on the predominance of characters from the respective language’s alphabet. After collecting the data, we found the texts contained various types of noise such as empty entries, incorrect sentences, duplicate entries, entries consisting solely of numbers, and other unwanted characters. These issues were systematically identified and removed to enhance the dataset’s quality. Additionally, unnecessary columns like "book" and "author" were deleted to focus exclusively on the translation pairs. We also removed religious Quranic verses from the dataset, as they were few in number and not dealt with correctly.
Alignment Verification:
As in the Rasaif websites—where we collected the translations from—the translations are created by human volunteers. Given the lack of detailed insights into their methods, and to ensure that each Arabic sentence was correctly aligned with its English translation, thereby maintaining the context and intended meaning, the authors manually verified the collected datasets. This verification process was crucial to confirm that the Arabic-English pairs were properly aligned and accurately conveyed the content of each other.
3.3 Comparative Analysis of ATHAR and Other Arabic Datasets
In this subsection, we analyze our dataset in comparison to existing classical and modern Arabic datasets, focusing on several linguistic measures: lexical diversity, stopword ratio, and the distribution of short versus long sentences, in addition to unique words count and dataset sizes.
For lexical diversity, we utilized the Measure of Textual Lexical Diversity (MTLD) index McCarthy (2005), which calculates the average length of text segments where the vocabulary diversity remains above a specified threshold—0.75 in our case—indicating how consistently a text maintains its lexical richness. The stopword ratio was calculated by determining the occurrence of stopwords relative to the total word count in the datasets. Short sentences were defined as any sentence containing 10 or fewer words, while long sentences are those with 30 or more words.
Before conducting the analysis, all datasets were standardized by removing redundant diacritics and letters. As detailed in Table 1, the ATHAR dataset boasts one of the highest MTLD scores, suggesting that the text can sustain a high level of lexical diversity over a large number of words. This implies that the vocabulary is varied and the text does not quickly repeat words. Furthermore, our dataset maintains a balanced representation of both short and long sentences, providing a stark contrast to the variable sentence lengths found in other datasets.
Dataset Attributes ATHAR Tanzil Arabic PCD KSUCCA OPUS-100-ar-en iwslt2017-ar-en multiun-ar-en
Dataset size 66K 187K 1.8M 1.9M 1M 241K 9.67M
Unique words count 138944 48104 720167 908771 370601 185390 841732
Lexical diversity (MTLD) 55.63 101.31 11.86 40.87 17.46 34.12 70.10
Ratio of stopwords (%) 26.04 30.35 24.62 24.71 27.59 29.67 21.31
Average length of sentences 20.78 34.35 9.26 25.33 8.39 13.86 22.89
Proportion of very short sentences (%) 24.06 11.18 76.57 41.28 79.81 45.98 23.07
Proportion of very long sentences (%) 23.11 47.53 0.00 24.44 4.71 7.04 26.61
Table 1: Overview of Linguistic Characteristics in Arabic Language Datasets: Size, Diversity, and Sentence Metrics
4 Evaluating State-of-the-Art LLMs on the ATHAR Dataset
In this section, we aim to evaluate the performance of state-of-the-art language models on classical Arabic translations using the ATHAR dataset. We selected four leading models for this analysis: GPT-4o, Llama-3 70B, Llama-3 8B, and Llama-2 7B.
Initially, we assessed the zero-shot capabilities of these models. Subsequently, we evaluated the Llama-3 8B and Llama-2 7B models under few-shot conditions. Finally, we focused on fine-tuning the Llama-3 8B model using two distinct methods: full fine-tuning, where all parameters of the model were adjusted, and LoRA Hu et al. (2021) fine-tuning, which only involved adjustments to a subset of newly added parameters.
The objective of these comprehensive experiments is to maximize the potential of these models, understand performance variations under different settings, and explore how the ATHAR dataset can bridge existing performance gaps.
In the following subsections, we will detail the hyperparameters and metrics used in our experiments and analyze the results.
4.1 Hyperparameters and Evaluation Metrics
Hyperparameters:
During inference, the generation decoding strategy involved setting the maximum number of new tokens to $2048$. Sampling strategies included Top-K and Top-P settings at $100$ and $0.95$, respectively, with a temperature parameter set at $0.3$.
For the fine-tuned models, specifically Llama-3 8B with both full and LoRA tuning, training was implemented in an instruction-input/response format. The input consisted of Arabic text, and the models were trained to generate the corresponding English translation as the response. The training dataset included $65k$ samples. Models were trained using $FP16$ precision, with a learning rate of $65e-$ 6, adjusted via a linear scheduler across three epochs. The batch size was set at $16k$ tokens, which was achieved by accumulating gradients of four samples twice, each sample having a maximum length of $2048$ tokens. An AdamW optimizer was utilized, with beta values of $0.90$ and $0.999$ for the first and second moment estimates, respectively.
Regarding the prompt structures used in our experiments, Table 3 details the specific prompt structures we utilized across zero-shot, few-shot, and fine-tuning settings.
Evaluation Metrics:
In assessing our models, we employed well-established metrics commonly used in translation evaluations: METEOR Banerjee and Lavie (2005), ROUGE-L Lin (2004), and SacreBLEU Post (2018). These metrics are all scored on a scale where higher values indicate better performance, though each has a different range. METEOR focuses on the alignment between the translation output and reference translations, considering synonymy and stemming. ROUGE-L measures the longest common subsequence, which is useful for evaluating the fluency of the text. SacreBLEU provides a consistent and comparable score across studies by standardizing the BLEU score calculation. Together, these metrics provide a comprehensive view of translation quality, covering aspects from accuracy to fluency.
4.2 Results and Discussion
Results:
The evaluation results, presented in Table 2, highlight significant variances in model performance across different settings. The GPT-4o model excelled in a zero-shot (ZS) setting, outperforming all other models with scores of $0.357$ on METEOR, $0.441$ on ROUGE-L, and $14.705$ on SacreBLEU. In contrast, the Llama-3 70B Instruct model, also evaluated in a zero-shot setting, registered slightly lower scores of $0.342$ on METEOR, $0.413$ on ROUGE-L, and $13.011$ on SacreBLEU. This disparity might reflect differences in training regimes or underlying model architectures.
In the same zero-shot context, both the Llama-3 8B Instruct and Llama-2 7B models displayed considerably lower performance, especially noted by sharp declines in their SacreBLEU scores to $0.339$ and $0.310$, respectively. These findings suggest inherent limitations in the zero-shot capabilities of these models for translation tasks.
Remarkable improvements were evident when the Llama-3 8B model was tested in a few-shot (FS) setting using just three samples. This adjustment led to marked score increases to $0.174$ on METEOR, $0.167$ on ROUGE-L, and $0.971$ on SacreBLEU, underscoring the potential of in-context learning to significantly boost model performance. In contrast, the few-shot enhancement for the Llama-2 7B model was minimal, which may indicate model-specific sensitivity to training data increments.
The Llama-3 8B model demonstrated further enhancements upon undergoing full fine-tuning, achieving a METEOR score of $0.275$, a ROUGE-L score of $0.336$, and a SacreBLEU score of $6.061$. Additionally, the LoRA tuning method, which involves less extensive modifications, also yielded improved results, with scores reaching $0.279$ on METEOR, $0.339$ on ROUGE-L, and $8.792$ on SacreBLEU.
Discussion:
The results presented in Table 2 underscore the challenges faced by state-of-the-art LLMs when tasked with translating Classical Arabic to English. By providing state-of-the-art models with targeted training opportunities, the ATHAR dataset not only boosts model performance but also contributes significantly to the broader NLP community’s understanding of and engagement with Classical Arabic. This dataset, therefore, holds substantial value, as it aids in developing more nuanced and capable translation systems.
Model METEOR $\uparrow$ ROUGE-L $\uparrow$ SacreBLEU $\uparrow$
GPT-4o + ZS 0.357 0.441 14.705
Llama-3 70B Instruct + ZS 0.342 0.413 13.011
Llama-3 8B Instruct + ZS 0.115 0.068 0.339
Llama-2 7B + ZS 0.116 0.099 0.310
Llama-3 8B Instruct + FS3 0.174 0.167 0.971
Llama-2 7B + FS3 0.036 0.039 0.408
Llama-3 8B + Full-Tuning 0.275 0.336 6.061
Llama-3 8B + LoRA 0.279 0.339 8.792
Table 2: Performance of State-of-the-Art LLMs on the Classical Arabic to English Translation Task. The table displays METEOR, ROUGE-L, and SacreBLEU scores for various models under different settings: zero-shot (ZS), few-shot with three samples (FS3), and fine-tuning (Full-Tuning & LoRA) on a 1k test set.
Model Prompt
[…]
Llama-3 8B Instruct + FS3 Llama-2 7B + FS3 Translate the following Classical Arabic text into English. Follow the provided examples for consistency and accuracy. Examples: Arabic: {arabtext}باجر قَالَ ابْن دُرَيْد وَهُوَ صَنَمٌ كَانَ لِلأَزْدِ فِي الْجَاهِلِيَّةِ وَمَنْ جَاوَرَهُمْ من طَيء وقضاعة. كَانُوا يَعْبُدُونَهُ. بِفَتْحِ الْجِيمِ, وَرُبَّمَا قَالُوا بَاجِرُ بِكَسْرِ الْجِيمِ English: According to Ibn Durayd, Bajar was an idol worshipped by the Azd tribe and neighboring tribes such as Tayyi’ and Quda’ah during the Jahiliyyah period. It is also pronounced Bajir. Arabic: {arabtext}(ذبل فِي آخر النُّسْخَة الَّتِي اعتمدتها فِي الطَّبْع) الْيَعْبُوبُ صَنَمٌ لِجَدِيلَةِ طَيْءٍ. وَكَانَ لَهُمْ صَنَمٌ أَخَذَتْهُ مِنْهُمْ بَنُو أَسَدٍ فَتَبَدَّلُوا الْيَعْبُوبَ بَعْدَهُ. قَالَ عبيد فَتَبَدَّلُوا الْيَعْبُوبَ بَعْدَ إِلهِهِمْ … صَنَمًا فَقَرُّوا يَا جديل واعذبوا (أَي لَا تَأْكُلُوا على ذَلِك وَلَا تشْربُوا) English: Al-Ya’bub was the idol of the Jadilah tribe of Tayyi’. They initially had a different idol, but after it was taken by Banu Asad, they adopted Al-Ya’bub in its place. ‘Abid said they replaced their former god with Al-Ya’bub, and as a result, O Jadilah, abstain from food and drink. Arabic: {arabtext}فَلم يزل الْفلس يعبد حَتَّى ظهرت دَعْوَة النَّبِي عَلَيْهِ السَّلَام فَبعث إِلَيْهِ على ابْن أَبِي طَالِبٍ فَهَدَمَهُ وَأَخَذَ سَيْفَيْنِ كَانَ الْحَارِثُ بن أبي شمرٍ الغساني, ملك غَسَّان قَلَّدَهُ إِيَّاهُمَا, يُقَالُ لَهُمَا مِخْذَمٌ وَرَسُوبٌ(وَهُمَا السَّيْفَانِ اللَّذَانِ ذَكَرَهُمَا عَلْقَمَةُ بْنُ عَبْدَةَ فِي شِعْرِهِ). فَقَدِمَ بِهِمَا عَلِيُّ بْنُ أَبِي طَالِبٍ عَلَى النَّبِيِّ صَلَّى اللَّهُ عَلَيْهِ وَسَلَّمَ فَتَقَلَّدَ أَحَدَهُمَا ثُمَّ دَفَعَهُ إِلَى عَلِيِّ بْنِ أَبِي طَالِبٍ, فَهُوَ سَيْفُهُ الَّذِي كَانَ يَتَقَلَّدُهُ English: Al-Fals continued to be worshipped until the advent of the Prophet. When ‘Ali ibn Abi Talib was sent to destroy it, he did so and took two swords, Mikhdham and Rasub, which had been given to Al-Fals by Al-Harith ibn Abi Shamir, the king of Ghassan. ‘Ali presented them to the Prophet, who kept one and returned the other to ‘Ali, who continued to wear it. Translate the following: Arabic: {Arabic text} English:
Llama-3 8B + Full-Tuning Llama-3 8B + LoRA Translate the following input text from Classical Arabic to English, please return only the translated text without any introductions or additions. ### Input: {Arabic text} ### Response:
Table 3: Prompt Structures Used and Their Corresponding Models in Zero-Shot (ZS), Few-Shot with Three Samples (FS3), and Full-Tuning Evaluation Experiments.
5 Conclusion
To conclude, we introduce the ATHAR dataset, which enhances the existing corpus of Classical Arabic datasets by incorporating a broader range of topics. Our evaluation of the current status of LLMs underscores the critical need for the ATHAR dataset within the fine-tuning and training pipelines. More broadly, this need highlights the necessity for more comprehensive Classical Arabic datasets to improve the quality of translation systems in this domain. Future work will aim to expand the ATHAR dataset to include an even wider array of texts and topics, thereby further enhancing translation quality.
Appendix A Preprocessing: Flipped Cells in Data Collection
During the scraping process, we encountered difficulties in extracting the English and Arabic texts from the containers (cells) because the Arabic texts were sometimes labeled as “flex-right” and English texts as “flex-left” in many instances, with the positions reversed in other cases. To address this, we counted the number of Arabic and English characters in each label and assigned the language based on the predominance of characters from either alphabet. Examples of such inconsistencies are provided below, where the labels for “flex-right” and “flex-left” are swapped, complicating the identification process:
Example.1 Arabic text on left and English text on right:
<div class="flex">
<div class="flex-right">
<span>Farewell my brother, whom it was my duty to help. The blessings and the mercy of God upon you</span>
</div>
<div class="flex-left">
{arabtext}
والسلام عليك أيها الأخ المفترض إسعافه ورحمة الله وبركاته
</div>
</div>
Example.2 Arabic text on right and English text on left:
<div class="flex">
<div class="flex-right">
{arabtext}
أنشدني أبو القاسم الصرويّ لنفسه" ويوم كيوم البين حرّا قطعته … على سابح طاوي الأياطل سابق، أخوض عليه جمرة القيظ حاسرا … كأنّي على الهجران في قلب عاشق
</div>
<div class="flex-left">
<span>Abu’l-Qasim Suri recited to me the following lines of his own: "O day as hot as day of lovers’ parting I spent upon a courser lean of flank! On him in summer’s wave like heart of lover burning with separation’s pain I sank."</span>
</div>
</div>